OpenStudy (anonymous):

find the zeros of the equation g(x)=x^2-25

OpenStudy (anonymous):

try factoring as the difference of two squares \[a^2-b^2=(a+b)(a-b)\]

OpenStudy (precal):

do as satellite73 suggested

OpenStudy (precal):

do you see what your b value is?

OpenStudy (anonymous):

Or, x^2=25.